A square plywood platform has a perimeter which is 9 times the length of a side, decreased by 15. Find the length of a side. Use the formula P = 4s.

Answer by CharlesG2(834)

A square plywood platform has a perimeter which is 9 times the length of a side, decreased by 15. Find the length of a side. Use the formula P = 4s.
P = 4s = 9s - 15
0 = 5s - 15
15 = 5s
3 = s
P = 4s = 4 * 3 = 12
P = 9s - 15 = 9 * 3 - 15 = 27 - 15 = 12